OpenStudy (anonymous):

I need to factor x^3+27

OpenStudy (anonymous):

Didnt someone explain this to you already?

OpenStudy (anonymous):

Idk, I typed it in wrong to begin with so I thought I would fix it

OpenStudy (anonymous):

okay well what do you notice about these two terms

OpenStudy (aroub):

just apply the rule

OpenStudy (anonymous):

27 has a cube root of 3

OpenStudy (anonymous):

Apply what rule?

OpenStudy (anonymous):

a^3+b^3=(a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

sum of cubes

OpenStudy (anonymous):

I don't know that rule... And so a=x and b=3?

OpenStudy (aroub):

ok the rule of sum of 2 cubes :x^3+b^3=(x+b) (x^2 -xb+b^2)

OpenStudy (anonymous):

yes

OpenStudy (aroub):

yes!

OpenStudy (anonymous):

YAY! Thank you! I have one more question.

OpenStudy (aroub):

okay! :)

OpenStudy (anonymous):

My final result is (x+3)(x^2-3x+9) is that right???

OpenStudy (aroub):

yes its true!!! :)

OpenStudy (anonymous):

YAY! Thank you!